# Diving Deep into Regions: Exploiting Regional Information Transformer for Single Image Deraining

### v1

@@ 4.2 Deraining Performance Evaluations
Table 1: Quantitative comparison (PSNR/SSIM) with state-of-the-art methods for single image rain streak removal. The best and second best results are colored with red and blue.
Datasets Rain200L Rain200H DID-Data DDN-Data SPA-Data
Metrics PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
DSC 27.16 0.8663 14.73 0.3815 24.24 0.8279 27.31 0.8373 34.95 0.9416
GMM 28.66 0.8652 14.50 0.4164 25.81 0.8344 27.55 0.8479 34.30 0.9428
DDN 34.68 0.9671 26.05 0.8056 30.97 0.9116 30.00 0.9041 36.16 0.9457
RESCAN 36.09 0.9697 26.75 0.8353 33.38 0.9417 31.94 0.9345 38.11 0.9707
PReNet 37.80 0.9814 29.04 0.8991 33.17 0.9481 32.60 0.9459 40.16 0.9816
MSPFN 38.58 0.9827 29.36 0.9034 33.72 0.9550 32.99 0.9333 43.43 0.9843
RCDNet 39.17 0.9885 30.24 0.9048 34.08 0.9532 33.04 0.9472 43.36 0.9831
MPRNet 39.47 0.9825 30.67 0.9110 33.99 0.9590 33.10 0.9347 43.64 0.9844
DualGCN 40.73 0.9886 31.15 0.9125 34.37 0.9620 33.01 0.9489 44.18 0.9902
SPDNet 40.50 0.9875 31.28 0.9207 34.57 0.9560 33.15 0.9457 43.20 0.9871
Uformer 40.20 0.9860 30.80 0.9105 35.02 0.9621 33.95 0.9545 46.13 0.9913
Restormer 40.99 0.9890 32.00 0.9329 35.29 0.9641 34.20 0.9571 47.98 0.9921
IDT 40.74 0.9884 32.10 0.9344 34.89 0.9623 33.84 0.9549 47.35 0.9930
DRSformer 41.23 0.9894 32.17 0.9326 35.35 0.9646 34.35 0.9588 48.54 0.9924
Regformer (ours) 41.51 0.9900 32.46 0.9353 35.43 0.9651 34.38 0.9591 48.60 0.9941
(1) Quantitative Results. The comparison results of our model and other related methods on different
@@
Eigen et al. Eigen et al. (2013) 21.31 0.7572
Pix2Pix Isola et al. (2017) 27.20 0.8359
AttentGAN Qian et al. (2018) 31.59 0.9170
Quan et al. Quan et al. (2019) 31.37 0.9183
IDT Xiao et al. (2022) 31.87 0.9313
Regformer (ours) 32.51 0.9725
@@
Table 2: Quantitative comparison with other related methods for single image raindrop removal on AGAN-Data Qian et al. (2018). The best and second best results are colored with red and blue.
datasets are shown in Table 1 and Table 2. Specifically, in Table 1, we show the results on rain streak datasets, while Table 2 shows our results on the raindrop dataset. For rain streak removal, we can clearly see that the PSNR and SSIM values of our method on each dataset exceed the existing solutions, or maintain a high competitive level. Especially on the Rain200L and Rain200H datasets, our method has greatly improved the restored results, and outperformed current methods by up to 0.26 dB and 0.28 dB. On other datasets, PSNR far exceeds the baseline Zamir et al. (2022), and at the same time refreshes the current SOTA level. This clearly demonstrates the effectiveness of our innovative approach to Regformer. In addition, it can be seen from the chart that the CNN-based method still maintains high competitiveness compared with the current Transformer-based method, but the performance of methods based on prior knowledge, such as DSC Luo et al. (2015) and GMM Li et al. (2016), is relatively backward. Although DRSformer also performs effective attention mechanism exploration based on Restormer, they lag behind our Regformer on both synthetic and real datasets since they do not explore the difference between different regions. In addition, for raindrop removal, we can clearly see that our method far exceeds current related methods by up to 0.64 dB, and the error rate decreased by nearly 60% (the ground truth’s SSIM is 1.0000).
@@
(2) Visualization Results. We present our rain streak removal results in detail in Figures 2 and 5. Notably, we present a visualization of our model’s ideas in Figure 1, 6, and 7, which focus on showing why and how our model works. Figures 2, 5 show the comparison results of ours and other existing representative methods on the real dataset SPA-Data, respectively. Clearly, we can see that in some areas, other compared methods do not remove the rain streak information well, or produce image distortion problems while removing the rain streak information. Different from other methods, our method preserves the original details of the image content area while removing the rain streaks, which fully demonstrates the superiority of our method. Due to space limitations, more visualization results will be shown in the appendix section.
4.3 Ablation Studies
[…]
(1) Effectiveness of Training Settings. Initially, we train our model using the default settings of
@@ 4.3 Ablation Studies
Table 3: Ablation study on Rain200L dataset. It reflects the role of each component in our Regformer.
Restormer (v1) and subsequently retrain it by adjusting some of its parameters (v2). The results can be seen in Table 3. For specific parameter details, please refer to the training details section. We fine-tune the number of layers and attention heads within each module to better adapt our model to the target task and establish a solid foundation for further modifications. The modifications in training settings serve to enhance our model’s performance, making it more effective in handling SID task.
@@
(2) Effectiveness of RTC and Region Mask Mechanism. Table 3 presents the results of incorporating the RTC and region mask mechanisms into our model. Clearly, adding the RTC module without the region mask mechanism does not significantly improve the model’s performance (compared to v2, v3 exhibits only a 0.01 dB increase). We add foreground mask (v6) and background mask (v7), respectively, and the performance improves by 0.20dB and 0.17dB, respectively. However, upon integrating the entire Region Mask mechanism, the network effectively captures feature information from different regions. This allows the model to efficiently comprehend the features of distinct regions and their fused attributes, thereby achieving a substantial performance enhancement (compared to v3, the performance of v4 improves by 0.17 dB; and relative to v5, the performance of Regformer is boosted by 0.25 dB). These results highlight the crucial role the region mask mechanism plays in Regformer, demonstrating its effectiveness in handling the single image deraining task.
[…]
5 Conclusion
@@
In this paper, we introduced Regformer, a novel SID model that combines the advantages of both the rain-affected and rain-unaffected regions. The Regformer model efficiently integrates the local and non-local features to effectively reconstruct high-quality images while preserving essential details in both the rain-affected and unaffected regions. Our proposed embedded region mask mechanism allows the model to selectively focus on crucial information from both regions. By employing rain masks and unaffected region masks, the model can better differentiate between the two regions and ultimately yield a feature map with region-specific features. Extensive experiments on rain streak and raindrop removal demonstrated the effectiveness of our model. The results showed that our model outperforms current methods in terms of both quantitative and qualitative evaluations. In conclusion, Regformer is a promising SID model, achieving superior performance by leveraging the power of Transformer and U-Net architectures while incorporating an effective region mask mechanism. Future work may involve exploring additional masking strategies for further improvement and extending the model to other image restoration tasks, including dehazing and denoising.
Acknowledgments
@@
This work is supported by the National Natural Science Foundation of China (62072151), Anhui Provincial Natural Science Fund for the Distinguished Young Scholars (2008085J30), Open Foundation of Yunnan Key Laboratory of Software Engineering (2023SE103), CCF-Baidu Open Fund and CAAI-Huawei MindSpore Open Fund.

### v2

@@ III-B2 Mixed Gate Forward Block
TABLE I: Quantitative comparison (PSNR/SSIM) with state-of-the-art methods for single image rain streak removal datasets, including Rain200L [40], Rain200H [40], DID-Data [41], DDN-Data [42], SPA-Data [19] and HQ-RAIN [43]. The best and second best results are colored with red and blue.
Datasets Venue Rain200L Rain200H DID-Data DDN-Data SPA-Data
Metrics - PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
DSC ICCV’15 27.16 0.8663 14.73 0.3815 24.24 0.8279 27.31 0.8373 34.95 0.9416
GMM CVPR’16 28.66 0.8652 14.50 0.4164 25.81 0.8344 27.55 0.8479 34.30 0.9428
DDN CVPR’17 34.68 0.9671 26.05 0.8056 30.97 0.9116 30.00 0.9041 36.16 0.9457
RESCAN ECCV’18 36.09 0.9697 26.75 0.8353 33.38 0.9417 31.94 0.9345 38.11 0.9707
PReNet CVPR’19 37.80 0.9814 29.04 0.8991 33.17 0.9481 32.60 0.9459 40.16 0.9816
MSPFN CVPR’20 38.58 0.9827 29.36 0.9034 33.72 0.9550 32.99 0.9333 43.43 0.9843
RCDNet CVPR’20 39.17 0.9885 30.24 0.9048 34.08 0.9532 33.04 0.9472 43.36 0.9831
MPRNet CVPR’21 39.47 0.9825 30.67 0.9110 33.99 0.9590 33.10 0.9347 43.64 0.9844
DualGCN AAAI’21 40.73 0.9886 31.15 0.9125 34.37 0.9620 33.01 0.9489 44.18 0.9902
SPDNet ICCV’21 40.50 0.9875 31.28 0.9207 34.57 0.9560 33.15 0.9457 43.20 0.9871
Uformer CVPR’22 40.20 0.9860 30.80 0.9105 35.02 0.9621 33.95 0.9545 46.13 0.9913
Restormer CVPR’22 40.99 0.9890 32.00 0.9329 35.29 0.9641 34.20 0.9571 47.98 0.9921
IDT TPAMI’22 40.74 0.9884 32.10 0.9344 34.89 0.9623 33.84 0.9549 47.35 0.9930
DRSformer CVPR’23 41.23 0.9894 32.17 0.9326 35.35 0.9646 34.35 0.9588 48.54 0.9924
HCN TPAMI’23 41.31 0.9892 31.34 0.9248 34.70 0.9613 33.42 0.9512 45.03 0.9907
Regformer - 41.51 0.9900 32.46 0.9353 35.43 0.9651 34.38 0.9591 48.60 0.9941
Regformer-L - 41.79 0.9901 32.69 0.9389 35.40 0.9444 34.40 0.9471 49.53 0.9941
Dataset HQ-RAIN
Methods LPNet PReNet JORDER-E RCDNet HINet SPDNet Uformer Restormer IDT DRSformer Regformer Regformer-L
PSNR 31.39 36.73 36.46 39.53 42.30 34.02 37.57 44.59 38.38 44.67 45.12 45.85
SSIM 0.9203 0.9657 0.9599 0.9771 0.9841 0.9359 0.9673 0.9905 0.9737 0.9898 0.9904 0.9920
LPIPS 0.1024 0.0350 0.0435 0.0242 0.0147 0.0813 0.0314 0.0076 0.0261 0.0081 0.0072 0.0059
Refer to caption rainy img_140 [19] Refer to caption Refer to caption GT (PSNR/SSIM) RESCAN (25.17/0.9084)[11] Refer to caption Refer to caption Restormer (21.56/0.9041) [38] DualGCN (23.82/0.9291) [44] Refer to caption Refer to caption SPDNet (23.51/0.9109) [45] IDT (20.90/0.9125) [17] Refer to caption Refer to caption DRSformer (19.38/0.8894) [15] Ours (33.96/0.9682)
LQ
SPDNet (24.73/0.8988)
Restormer (20.23/0.8619)
DualGCN (26.40/0.9507)
GT (PSNR/SSIM)
IDT (23.88/0.9279)
DRSformer (19.34/0.8539)
Ours (36.00/0.9762)
Figure 4: Visual comparison of each method on the SPA-Data [19] dataset. The visualization illustrates the performance of various methods in processing rain streak-like content features. Our model stands out by preserving more precise details while effectively removing rain streaks, demonstrating its superiority over previous methods.
Refer to caption rainy img_465 [19] Refer to caption Refer to caption GT (PSNR/SSIM) RESCAN (24.04/0.9024) [11] Refer to caption Refer to caption Restormer (23.48/0.9102) [38] DualGCN (22.09/0.8831) [44] Refer to caption Refer to caption SPDNet (20.46/0.0.8593)[45] IDT (27.60/0.9273)[17] Refer to caption Refer to caption DRSformer (23.79/0.9140) [15] Ours (38.26/0.9775)
Figure 5: Visual comparison on the SPA-Data [19] dataset. To facilitate the display of the results, the local graph has been rotated 90 degrees clockwise. Compared to previous solutions, our method excels at distinguishing content-like rain streak noise and preserving the original details of the input image while effectively removing rain streaks, demonstrating its superiority over previous methods.
Refer to caption rainy img_124 Refer to caption Refer to caption Refer to caption Refer to caption GT (PSNR/SSIM) DDN (26.09/0.7232) Restormer (29.72/0.8532) DualGCN (27.87/0.8330) Refer to caption Refer to caption Refer to caption Refer to caption SPDNet (29.72/0.8532) IDT (28.21/0.8535) DRSformer (31.00/0.8966) Ours (31.69/0.9024)
Figure 6: Visual comparison of different methods on Rain200H [40] dataset. Clearly, our Regformer model can perform more accurately in detail and texture recovery than other approaches.
IV Experiments
IV-A Experimental Settings
Evaluated Datasets. We evaluate the performance of our model on several public benchmarks, including Rain200L [40], Rain200H [40], DID-Data [41], DDN-Data [42], SPA-Data [19] and HQ-RAIN [43]. Specifically, Rain200L and Rain200H consist of 1,800 synthetic rainy image pairs for training and 200 synthetic test image pairs. DID-Data and DDN-Data contain 12,000 and 12,600 synthetic rainy image pairs for training, respectively, along with 1,200 and 1,400 rainy image pairs for testing. SPA-Data is a real-world dataset that includes 638,492 training image pairs and 1,000 testing image pairs. Meanwhile, the training and testing set of the HQ-RAIN dataset contains 4,500 and 500 synthetic images, respectively. We further evaluate Regformer on a real-world raindrop dataset AGAN-Data [13], which includes 1,110 image pairs.
Compared Methods. For rain streak removal, we compare our Regformer model with a variety of state-of-the-art methods, including prior knowledge-based approaches like DSC [46] and GMM [47], CNN-based methods such as DDN [42], RESCAN [11], LPNet [48], JORDER-E [49], PReNet [50], MSPFN [51], RCDNet [52], HINet [53], MPRNet [54], DualGCN [44], SPDNet [45] and HCN [55]. In addition, we consider Transformer-based methods, including Uformer [39], Restormer [38], IDT [17] and DRSformer [15]. We refer to the experimental data provided by DRSformer and adhere to their original experimental settings for further analysis. Our experimental results are presented in Table I. In addition, we also test our model on one real-world raindrop dataset AGAN-Data [13] to measure our model performance more broadly. We compare our model with five methods, including Eigen et al. [56], Pix2Pix [57], AttentGAN [13], Quan et al. [26] and IDT [17]. Our experimental results are presented in Table II. We employ PSNR (Peak Signal-to-Noise Ratio) [58] and SSIM (Structural Similarity Index Measure) [59] as the primary evaluation metrics. Note that we calculate the metrics on the Y channel of YCbCr like the previous works [15, 52, 40].
Implementation and Training Details. In our Regformer model, we configure the layers L0, L1, L2 and L3 to 4, 4, 4, 4, and the corresponding attention heads to 6, 6, 6, 6. we set the initial channel to 48. For Regformer-L model, we configure the layers to 6, 6, 6, 6 and initial channel to 60 insteadily. During training, we employ the AdamW [60] optimizer with parameters $\beta 1$ = 0.9, $\beta 2$ = 0.999, and weight decay = 1e-4. In the experiment part, we set $n$ = 2, $k_{1}$ = 3 and $k_{2}$ = 5 in the MGFB. Please note that in different tasks or different datasets, we may modify these hyperparameters to achieve better results. Concurrently, we follow DRSformer’s settings [15], which utilize the L1 loss as the loss function. We set the initial learning rate to 3e-4 and apply a cosine annealing algorithm to gradually decay it to 1e-6. The patch size is set to $128\times 128$, and we perform a total of 300K iterations during training for Regformer, while for Regformer-L we use totally 400K iterations since it needs more time to converge. We use two NVIDIA GeForce RTX 3090 GPUs with 24GB memory each for training.
IV-B Deraining Performance Evaluations
(1) Quantitative Results. The comparison results of our model and other related methods on different datasets are shown in Table I and Table II. Specifically, in Table I, we show the results on rain streak datasets, while Table II shows our results on the raindrop dataset. For rain streak removal, we can clearly see that the PSNR and SSIM values of our method on each dataset exceed the existing solutions, or maintain a high competitive level. Especially on the Rain200L and Rain200H datasets, our method has greatly improved the restored results, and outperformed current methods by up to 0.26 dB and 0.28 dB. On other datasets, PSNR far exceeds the baseline [38], and at the same time refreshes the current SOTA level. Moreover, the performance of Regformer-L outperforms DRSformer [15] by up to 1.18dB. This clearly demonstrates the effectiveness of our innovative approach to Regformer.
In addition, it can be seen from the chart that the CNN-based method still maintains high competitiveness compared with the current Transformer-based method, but the performance of methods based on prior knowledge, such as DSC [46] and GMM [47], is relatively backward. Although DRSformer also performs effective attention mechanism exploration based on Restormer, they lag behind our Regformer on both synthetic and real datasets since they do not explore the difference between different regions. In addition, for raindrop removal, we can clearly see that our method far exceeds current related methods by up to 0.64 dB.
@@ IV-B Deraining Performance Evaluations
Eigen et al. [56] 21.31 0.7572
Pix2Pix [57] 27.20 0.8359
AttentGAN [13] 31.59 0.9170
Quan et al. [26] 31.37 0.9183
IDT [17] 31.87 0.9313
Regformer (ours) 32.51 0.9420
@@
TABLE II: Quantitative comparison with other related methods for single image raindrop removal on AGAN-Data [13]. The best and second best results are colored with red and blue.
@@
(2) Visualization Results. We present our rain streak removal results in detail in Figures 4, 5 and 6. Notably, we present a visualization of our model’s concepts in Figures 1, 8, 9, and 10, which focus on explaining why and how our model works. Figure 4 shows the results of different methods processing rain streak-like content features, while Figure 5 shows the results of different methods processing content-like rain streak noise. Clearly, in some areas, other compared methods do not remove the rain streak information well or produce image distortion problems while removing the rain streak information. In contrast, our method better distinguishes the two types of features and accurately restores high-quality results, preserving the original details of the image content area while effectively removing the rain streaks. Further more, in Figure 6, our method clearly restore images with better texture recovery quality. Above visualization results fully demonstrate the superiority of our method.
Figure 7: Performance line chart and training process line chart on Rain200L [40] dataset.
IV-C Ablation Studies
Input
Rain Mask
Figure 8: These mask feature maps on the Rain200L dataset [40] represent the masks as they are processed within the network during inference. To effectively illustrate our high-quality mask, we have averaged the masks across their respective channels and transformed them into RGB images for a more direct and visual representation.
Input
Unaffected Mask
img_1
img_2
img_3
img_4
img_5
img_6
Figure 9: The visualization image results of unaffected region mask we generated on the DDN dataset [42]. They are one image with rain content added in different directions and varying levels. As can be seen in the figure, most of the rain region and the unaffected region has been successfully distinguished.
In the ablation study, we utilize the Rain200L dataset to train and evaluate our network. We take Restormer [38] as the Baseline and modify some of its parameters to enhance its performance. In Table III, Baseline1 (v1) denotes the initial Restormer; baseline2 (v2) denotes the Restormer with optimized parameters; w/RTC1 (v3) refers to the addition of RTC without employing the RegionMask mechanism (i.e., no mask operation); and w/RTC2 (v4) signifies the incorporation of the RegionMask mechanism. The term w/MGFB (v5) represents the integration of the MGFB. Moreover, we show the performance line chart and training process line chart in figure 7 to evaluate our model performance more comprehensively. This ablation study aims to provide a clear understanding of the individual component’s contributions and the overall effectiveness of our proposed Regformer model.
Figure 10: Feature map after unaffected region masked attention mechanism obtained during inference on Rain200H [40] dataset. It can be clearly seen that the background content feature can be successfully distinguished.
(1) Effectiveness of Training Settings. Initially, we train our model using the default settings of Restormer (v1) and subsequently retrain it by adjusting some of its parameters (v2). The results can be seen in Table III. For specific parameter details, please refer to the training details section. We fine-tune the number of layers and attention heads within each module to better adapt our model to the target task and establish a solid foundation for further modifications. The modifications in training settings serve to enhance our model’s performance, making it more effective in handling SID task.
(2) Effectiveness of RTC and Region Mask Mechanism. Table III presents
TABLE III: Ablation study on Rain200L dataset. It reflects the role of each component in our Regformer.
@@ IV-C Ablation Studies
the results of incorporating the RTC and region mask mechanisms into our model. Clearly, adding the RTC module without the region mask mechanism does not significantly improve the model’s performance (compared to v2, v3 exhibits only a 0.01 dB increase). We add foreground mask (v6) and background mask (v7), respectively, and the performance improves by 0.20dB and 0.17dB, respectively. However, upon integrating the entire Region Mask mechanism, the network effectively captures feature information from different regions. This allows the model to efficiently comprehend the features of distinct regions and their fused attributes, thereby achieving a substantial performance enhancement (compared to v3, the performance of v4 improves by 0.17 dB; and relative to v5, the performance of Regformer is boosted by 0.25 dB). These results highlight the crucial role the region mask mechanism plays in Regformer, demonstrating its effectiveness in handling the single image deraining task.
(3) Effectiveness of MGFB. Table III illustrates the impact of incorporating MGFB into our model. As evident from the table, adding MGFB without introducing RTC (v5) does not significantly enhance the model’s restoration performance (compared to v2, the PSNR of v5 increases by a mere 0.02 dB). This outcome can be attributed to MGFB being primarily designed as a mixed-scale feature extraction mechanism for different Region Mask attention mechanisms, intended to thoroughly extract features across various scales and receptive fields in distinct regions. Without employing the Region Mask attention mechanism, MGFB is restricted to extracting information from a single-layer region, which limits its performance. However, when RTC2 is integrated, we see a notable performance improvement (Regformer’s PSNR is enhanced by 0.09 dB compared to v4), which directly substantiates the validity and rationality of our approach. This highlights the effectiveness of MGFB within our Regformer for SID.
V Conclusion
@@
In this paper, we introduced Regformer, a novel SID model that combines the advantages of both the rain-affected and rain-unaffected regions. The Regformer model efficiently integrates the local and non-local features to effectively reconstruct high-quality images while preserving essential details in both the rain-affected and unaffected regions. Our proposed embedded region mask mechanism allows the model to selectively focus on crucial information from both regions. By employing rain masks and unaffected region masks, the model can better differentiate between the two regions and ultimately yield a feature map with region-specific features. Extensive experiments on rain streak and raindrop removal demonstrated the effectiveness of our model. The results showed that our model outperforms current methods in terms of both quantitative and qualitative evaluations. In conclusion, Regformer is a promising SID model, achieving superior performance by leveraging the power of Transformer and U-Net architectures while incorporating an effective region mask mechanism.
Acknowledgment
@@
This work is supported by the National Natural Science Foundation of China (62072151), Anhui Provincial Natural Science Fund for the Distinguished Young Scholars (2008085J30), Open Foundation of Yunnan Key Laboratory of Software Engineering (2023SE103), CCF-Baidu Open Fund, CAAI-Huawei MindSpore Open Fund and the Fundamental Research Funds for the Central Universities (Grant No. PA2024GDSK0096).